OpenStudy (anonymous):

Solve the system, using substitution. 4x + y = 4 and y = 4 - 4x A. no solution B. infinitely many solutions I put A is this correct?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

No, that is incorrect. The answer is B.

OpenStudy (anonymous):

Im getting mixed signals...

OpenStudy (anonymous):

there are infinite solutions

OpenStudy (anonymous):

lol. since 4=4 i don't think there is infinite solutions

OpenStudy (anonymous):

$$(x,y) = (1, 0)$$ is an answer.

OpenStudy (anonymous):

as is $$(x,y)=(1/2, 2)$$.

OpenStudy (anonymous):

nomothetis how did you get this ?

OpenStudy (anonymous):

I am getting a little MORE confused... ^_^ lol so many diff answers!!

OpenStudy (anonymous):

just try some (x,y) = (0.5, 2) is another solution

OpenStudy (anonymous):

:P sorry i am retaking stuff from like the beginning of the year :P So then do u all agree with jimmy?

OpenStudy (anonymous):

Here's how you prove that you have an infinite number of solutions: Can you make get equation 2 by multiplying equation 1 by something? Let's consider it: $$4x + y = 4$$ This is the same thing as: $$y = 4 - 4x$$ Since all we need to do is pass the x to the other side. So the two equations are really the same equation. So any points that is a solution for the first equation will work for the second equation. But if you look at the first equation, you see that it's the equation for a line. Therefore, any point on that line will be a solution to the second equation as well. This means that you have an infinite number of solutions.

OpenStudy (anonymous):

oh nomo siad that one try (x,y) = (0.25,3) then!!!

OpenStudy (anonymous):

Ok so its B?!?!?! are we positive on this answer now? lol

OpenStudy (anonymous):

i don't agree since it clearly states that "solve by substitution"

OpenStudy (watchmath):

B) is correct, even if you use substitution. The substitution gives 4=4 which means that the statement is always true for any (x,y) as long as (x,y) is on the line 4x+y=4.

OpenStudy (anonymous):

right nomo - the 2 equations are the same and u can fit infinite values of x and corresponding values of y into the equation

OpenStudy (anonymous):

ok i see. that makes sense watchmath

OpenStudy (anonymous):

absolutely positively b !! lol

OpenStudy (anonymous):

Okay!! haha finally!! (: lol